There are at most 12 bicycles and tricycles in a school playground. There are at least 17 wheels altogether. Let b equal the number of bicycles and t equal the number of tricycles. Which system describes this situation?




b + t <= 12 2b + 3t >=17

b + t <=12 2b + 3t <= 17

b + t <= 12 2b + 3t > 17

b + t < 12 2b + 3t >=17

Found 2 solutions by addingup, stanbon:
Answer by addingup(3677)

There are at most 12 means there are 12 or less: b+t<=12
At least 17 means there are 17 or more: 2b+3t>= 17
Thus, b + t <= 12; 2b + 3t >=17 is your answer

Answer by stanbon(75887)

There are at most 12 bicycles and tricycles in a school playground. There are at least 17 wheels altogether. Let b equal the number of bicycles and t equal the number of tricycles. Which system describes this situation?
----
b + t <= 12
2b + 3t >= 17
Cheers,
Stan H.
----------------

b + t <= 12 2b + 3t >=17
b + t <=12 2b + 3t <= 17
b + t <= 12 2b + 3t > 17
b + t < 12 2b + 3t >=17